Given function h(×)=2×+9 determine f(2)+3
1
Expert's answer
2021-09-21T05:02:37-0400

"f(x) = 2x+9\\
f(2) = 2(2) + 9 = 13\\
f(2) +3 = 2(2) +9+3 = 16"